A water tank can be filled in 20 hours by an inlet pipe and emptied in 24 hours by an outlet pipe. How long will it take to fill the tank if both pipes are left open?
Answer by sudhanshu_kmr(1152)


portion of tank will fill in one hour = 1/20 -1/24 = 1/120

so, it will take 120 hours to fill the tank of both pipes are open.